could you please help me solve for x in this equation:
2x+y=5
x-y=1
could you by any chance please explain in the easiest way possible i tried the solver but i wasnt understanding the solution very well so if you could help me out step by step that would be great thank you.

Answer by Abbey(339)

Because x-y=1, we can add y to both sides of the equation and get
x=1+y
Using 2(x)+y=5:
substitute the (1+y)in for x and write the equation:
2(1+y)+y=5
clear the parentheses:
2+2y+y=5
2+3y=5
subtract 2 from both sides:
3y=3
divide both sides by 3
y=1
use x-y=1 to find x
x-(1)=1
x=2
the solution is (2,1)